-4y-11x=36
20=-10x-10y

Answer by mananth(16946)


-11.00 x -4.00 y = 36.00 .............1

10.00 x + 10.00 y = -20.00 .............2
Eliminate y
multiply (1)by 5.00
Multiply (2) by 2.00
-55.00 x \ -20.00 y = 180.00
20.00 x + 20.00 y = -40.00
Add the two equations
-35.00 x = 140.00
/ -35.00
x = -4.00
plug value of x in (1)
-11.00 x -4.00 y = 36.00
44.00 -4.00 y = 36.00
-4.00 y = 36.00 -44
-4.00 y = -8.00
y = 2.00
x=-4.00
y=2.00
m.ananth@hotmail.ca